OpenStudy (nickymail48):

(x^2+4x+2)-(5x^2-5x+6)

OpenStudy (adamaero):

terms raised to the same power can be added/subracted

OpenStudy (anonymous):

(x3 + 3x2 + 5x – 4) – (3x3 – 8x2 – 5x + 6) = (x3 + 3x2 + 5x – 4) – 1(3x3 – 8x2 – 5x + 6) = (x3 + 3x2 + 5x – 4) – 1(3x3) – 1 (–8x2) – 1(–5x) – 1(6) = x3 + 3x2 + 5x – 4 – 3x3 + 8x2 + 5x – 6 = x3 – 3x3 + 3x2 + 8x2 + 5x + 5x – 4 – 6 = –2x3 + 11x2 + 10x –10

OpenStudy (adamaero):

that is a different question